OpenStudy (anonymous):

Please help me! "Determine the acceleration (in m/s/s) of a rider on the Cajun Cliffhanger (a barrel ride at an amusement park) if the rider makes 8.44 revolutions around the 5.23-m diameter circle in 28.8 seconds."

OpenStudy (anonymous):

You need to know these formulas. \[\omega=\frac{ \theta }{ t } , v= r \omega, a=\frac{ v ^{2} }{ r }\] You also need to know that: 1) r= radius (you are given the diameter) 2) how to convert from \[rev \to \theta \] . . . . \[\theta =rev * \frac{ 2\pi} { 1 }\] That should allow you to solve the problem. I hope this is right, I'm a little rusty.

OpenStudy (anonymous):

Yea that'll do it.

OpenStudy (anonymous):

The period \[T=2\pi radians\] \[\frac{ 8.44revs }{ 28.8\sec }*\frac{ 2\pi radians }{ 1rev }=\frac{ 211\pi radians }{ 360 \sec}=\omega\]

OpenStudy (anonymous):

Change rotational velocity into translational velocity. \[v=r \omega \] \[v=\left( \frac{ 5.23m }{ 2 } \right)\left( \frac{ 211\pi rads }{ 360\sec } \right)\approx4.815\frac{ m }{ \sec }\]

OpenStudy (anonymous):

Then the centripetal acceleration. We know it's centripetal because it's going around in a circle. \[\alpha=\frac{ v^2 }{ r }=\frac{ 4.815 \frac{ m }{ s }}{ 5.23m*.5 }\approx1.84\frac{ m }{ s^2 }\]

OpenStudy (anonymous):

awesome!!!! thank you guys so much!!!